# Physics-Constrained Fine-Tuning of Flow-Matching Models for Generation and Inverse Problems

### v1

@@ Appendix C Method: Implementation Details
The memoryless noise schedule $\sigma(t)=\sqrt{2\eta_{t}}$ in our setting of $\beta_{t}=t$ and $\gamma_{t}=1-t$ can be simplified to
$\sigma(t)=\sqrt{\frac{2(1-t)}{t}},$
which causes numerical problems for $t=0$. Furthermore, it forces the control $u$ to be close to zero for $t$ close to $1$. Following the original paper, we instead use
$\sigma(t)=\sqrt{\frac{2(1-t+h)}{t+h}},$
@@
Our PDE based fine-tuning residuals are particularly sensitive to noise. Therefore, we must ensure that sampled trajectories are fully denoised. When using memoryless noise, we found that we had to increase the number of steps per sampling trajectory to achieve the same quality of output compared to the ODE-based sampler. For denoising, especially the last steps of sample generation are important. As a trade-off between computational cost and accuracy, we implemented the following strategy: before fine-tuning we decide on a coarse time grid
$0=t_{0}<\dots<t_{T-1}=1$
that will also be used for all test-time rollouts, ensuring a fair comparison with the base model. During fine-tuning we subdivide the last $K_{\text{sub}}$ intervals geometrically: the last interval $t_{T-2},t_{T-1}]$ is split into $K_{\text{sub}}$ equal sub-steps, the previous one into $K_{\text{sub}}\!-\!1$, and so on. Formally, interval $t_{T-m-1},t_{T-m}]$ receives $K_{\text{sub}}-m$ internal points for $m=0,\dots,K_{\text{sub}}-1$
At inference time we revert to the original $T$-point grid for both base and fine-tuned models. Thus every quantitative comparison reported in the main paper is based on identical rollout budgets, while the fine-tuned model benefits from the cleaner gradients provided by the adaptive tail refinement during training.
@@
Full Training Algorithm
@@
Algorithm 1 details the complete optimisation loop used in all experiments. Starting from the pre-trained base flow $v^{\text{base}}$ we attach two residual heads that (i) condition the state flow $v_{t,x}^{\text{ft}}$ on the latent parameter $\alpha$ and (ii) predict the parameter flow $v_{t,\alpha}^{\text{ft}}$. The inverse predictor $\varphi$ is first pre-trained on base samples and then frozen, providing surrogate target flows for $\alpha$. Each epoch rolls both the base and fine-tuned trajectories on an augmented time grid, solves the lean-adjoint equation, and updates only the fine-tune parameters $\theta$ through the Adjoint-Matching loss.
@@
Algorithm 1 Adjoint Matching on Joint Evolution
Input: initialise core $v^{\text{ft}}\leftarrow v^{\text{base}}$, add residual-style output heads to get $v^{\text{ft}}_{x}$, $v^{\text{ft}}_{\alpha}$
@@
1:Pretrain $\varphi$ based on $x_{1}$ samples generated with $v^{\text{base}}_{x}$ by minimizing $\mathcal{R}_{\text{weak}}(x_{1},\varphi(x_{1}))$
2:Freeze weights of $v^{\text{base}}_{x}$ and $\varphi$, denote trainable parameters as $\theta$
3:for number of epochs do
4:  $x_{0}\sim\mathcal{N}(0,I),\quad\alpha_{0}\sim\mathcal{N}(0,I)$
5:  $\bar{T}\leftarrow\textsc{get\_augmented\_time}(T)$
6:  for $i\in\lvert\bar{T}\rvert$ do
7:   $t\leftarrow\bar{T}_{i},\quad t^{+}\leftarrow\bar{T}_{i+1},\quad h\leftarrow t^{+}-t$
8:   $v_{x,t}^{\text{base}}\leftarrow v^{\text{base}}(x_{t},t)$
9:   $v_{\alpha,t}^{\text{base}}\leftarrow\frac{\hat{\alpha}_{1}^{\text{ft}}-\alpha_{t}}{1-t}$
10:   $v_{x,t}^{\text{ft}},\ v_{\alpha,t}^{\text{ft}}\leftarrow v^{\text{ft}}(x_{t},\alpha_{t},v_{\alpha,t}^{\text{base}},t)$
11:   $x_{t^{+}}^{\text{base}}\leftarrow\textsc{sde\_step}(x_{t}^{\text{base}},v_{x,t}^{\text{base}},\sigma(t),h)$
12:   $x_{t^{+}}^{\text{ft}}\leftarrow\textsc{sde\_step}(x_{t}^{\text{ft}},v_{x,t}^{\text{ft}},\sigma(t),h)$
13:   $\alpha_{t^{+}}^{\text{ft}}\leftarrow\textsc{sde\_step}(\alpha_{t}^{\text{ft}},v_{\alpha,t}^{\text{ft}},\sigma(t),h)$
14:  end for
15:  $\tilde{a}\leftarrow\textsc{solve\_lean\_adjoint}(x^{\text{base}},x^{\text{ft}},\alpha^{\text{ft}})$
16:  $)\theta\leftarrow\textsc{gradient\_descent}(\theta,\,\mathcal{L}(\tilde{u};\tilde{X},\tilde{a}$))
17:end for
18:return $v^{\text{ft}}$, $\varphi$
Appendix D Details: Experimental Results
Hyperparameters for Experimental Results
Before starting fine-tuning, we train the inverse predictor $\varphi$. In 2D, we generate 1024 samples $x_{1}$ from the base model, in 3D we use 512. On these samples, the inverse predictor is trained for 100 epochs to minimize the weak residual $\mathcal{R}_{\text{weak}}(x_{1},\varphi(x_{1}))$. As the optimizer, we use Adam with a learning rate of $1\times 10^{-4}$ and a batch size of 8.
For fine-tuning, all runs use Adam with a fixed learning rate $2\times 10^{-5}$ and are trained on a single NVIDIA L40S GPU (48 GB). The settings that differ across tasks are summarised below.
Darcy Experiments
Denoising Misspec.
Time horizon $T$ 100 100
Tail fraction $K_{\!{\rm last}}$ 0.20 0.20
Evaluation steps $K$ 20 20
Tail refinement $K_{\!{\rm sub}}$ 5 5
# test functions 4096 (one per grid point)
$\sigma_{\min},\sigma_{\max}] [$3,10][3,10\,] [1,20][1,20\,]
Epochs 200 1000
Batch size 6 6
Runtime $\approx$30 min $\approx$2.5 h
In 2D, it is computationally feasible to sample one test function centered at each grid point. To still have stochasticity in location of test functions but ensure good spatial coverage, we sample small offsets that we add to each grid point. The fine-tuning network has $\approx 19$ M parameters and the inverse predictor $\approx 9$ M.
Acoustic Experiment
Sampling steps $T$ 40
Tail fraction $K_{\!{\rm last}}$ 0.25
Evaluation steps $K$ 10
Tail refinement $K_{\!{\rm sub}}$ 3
# test functions 1000
Super-resolution factor 4
$\sigma$ range [1,20][1,20\,]
Epochs 400

### v2

@@ D.4 Adjoint Matching Details
Domingo-Enrich et al. (2025) define a generative process with noise schedule $\sigma^{2}(t)$ to be memoryless, if and only if $\sigma^{2}(t)=2\eta_{t}+\chi(t)$ with $\chi:[0,1]\to\mathbb{R}$ chosen such that
$\forall t\in(0,1)\quad\lim_{t^{\prime}\to 0}\beta_{t^{\prime}}\ \exp\Biggl(-\int_{t^{\prime}}^{t}\frac{\chi(s)}{2\gamma^{2}_{s}}\ ds\Biggr)=0.$
Specifically, they refer to $\sigma(t)=\sqrt{2\eta_{t}}$ as the memoryless noise schedule. In our setting of $\beta_{t}=t$ and $\gamma_{t}=1-t$, the memoryless noise schedule can be simplified to
$\sigma(t)=\sqrt{\frac{2(1-t)}{t}}.$
Lemma 1 (Scaling of memoryless noise).
Consider a generative process as in 1 with $\beta_{t}=t$ and $\gamma_{t}=1-t$. For $0\leq\kappa<1$, the schedule $\sigma^{2}(t)=(1-\kappa)\ 2\eta_{t}$ is memoryless.
Proof.
First, we consider the integral term. The desired $\sigma^{2}(t)=(1-\kappa)\ 2\eta_{t}$ implies that
$\chi(t)=-2\kappa\eta_{t}=-2\gamma_{t}^{2}\ \frac{\kappa}{t(1-t)}.$
With this, we can simplify the integral term:
$-\int_{t^{\prime}}^{t}\frac{\chi(s)}{2\gamma^{2}_{s}}\ ds=\kappa\int_{t^{\prime}}^{t}\frac{1}{s(1-s)}\ ds$ $=\kappa\int_{t^{\prime}}^{t}\frac{1}{s}+\frac{1}{1-s}\ ds$
$=\kappa\ \bigl[\log(s)-\log(1-s)\bigr]_{t^{\prime}}^{t}$
$=\kappa\Biggl(\log\frac{t}{1-t}-\log\frac{t^{\prime}}{1-t^{\prime}}\Biggr).$
Thus, for an arbitrary fixed $t\in(0,\,1)$, it holds that
$\beta_{t^{\prime}}\,\exp\Biggl(-\int_{t^{\prime}}^{t}\frac{\chi(s)}{2\gamma^{2}_{s}}\ ds\Biggr)$ $=t^{\prime}\,\exp\Biggl(\kappa\ \Bigl(\log\frac{t}{1-t}-\log\frac{t^{\prime}}{1-t^{\prime}}\Bigr)\Biggr)$
$=t^{\prime}\,\left(\frac{t}{1-t}\right)^{\kappa}\,\left(\frac{t^{\prime}}{1-t^{\prime}}\right)^{-\kappa}$
[…]
Therefore, scaling down the noise schedule by a factor $1-\kappa$ is justified and consistent with the theory provided in Domingo-Enrich et al. (2025).
Heuristics.
Still, the term $\eta_{t}$ causes numerical problems for $t=0$. Furthermore, it forces the control $u$ to be close to zero for $t$ close to $1$. Following the original paper, we instead use
$\eta_{t}=\frac{1-t+h}{t+h},$
@@
While $\kappa$ is an effective tool to mitigate residual noise in final samples, increasing slows down training. As another way of improving sample quality without adding computational cost, we consider non-uniform time grids when sampling memoryless rollouts. We observe that the most critical phases of sampling are at the beginning, where noise magnitudes are the highest, and at the end, where final denoising happens. Therefore, we tilt the uniform grid towards the endpoints:
Let $S\in\mathbb{N}$ be the number of steps and define the uniform grid $t_{i}=i/S$ for $i=0,\dots,S$. We tilt this grid toward the endpoints by first applying a cosine–ease mapping
$g(t)\;=\;\tfrac{1}{2}\,\bigl(1-\cos(\pi t)\bigr),\qquad t\in[0,1],$
which has $g(0)=0$, $g(1)=1$. For a mixing parameter $q\in[0,1]$, the tilted times are the convex combination
$\tilde{t}_{i}\;=\;(1-q)\,t_{i}\;+\;q\,g(t_{i}),\qquad i=0,\dots,S.$
$\{\tilde{t}_{i}\}$ is strictly increasing and distributes grid points more densely near $t=0$ and $t=1$ for $q>0$. For PDE experiments, we use $q=0.9$. This heuristic was not needed for latent-space models.
@@
D.5 Full Training Algorithm
@@
Algorithm 1 details the complete optimisation loop used in all experiments. Starting from the pre-trained base flow $v^{\text{base}}$ we attach two residual heads that (i) condition the state flow $v_{t,x}^{\text{ft}}$ on the latent parameter $\alpha$ and (ii) predict the parameter flow $v_{t,\alpha}^{\text{ft}}$. Their respective output layers are initialized to zero. The inverse predictor $\varphi$ is first pre-trained on base samples and then frozen, providing surrogate target flows for $\alpha$. Each epoch rolls both the base and fine-tuned trajectories on the tilted time grid, solves the lean-adjoint equation, and updates only the fine-tune parameters $\theta$ through the Adjoint-Matching loss.
@@
Algorithm 1 Adjoint Matching with Joint State–Parameter Evolution
Notation: $v_{x}^{\mathrm{base}},v_{x}^{\mathrm{ft}},v_{\alpha}^{\mathrm{ft}}$ denote functions (neural vector fields). Evaluations at time $t$ are written in bold, e.g. $\mathbf{v}_{x,t}^{\mathrm{base}}:=v_{x}^{\mathrm{base}}(x_{t},t)$.
Input: base state field $v_{x}^{\mathrm{base}}$; initialize $v^{\mathrm{ft}}\leftarrow v_{x}^{\mathrm{base}}$ and add heads for $v_{x}^{\mathrm{ft}}$, $v_{\alpha}^{\mathrm{ft}}$.
@@
1:Pretrain $\varphi$ on $x_{1}$ samples from $v_{x}^{\mathrm{base}}$ by minimizing $\mathcal{R}_{\mathrm{weak}}(x_{1},\varphi(x_{1}))$.
2:Freeze $v_{x}^{\mathrm{base}}$ and $\varphi$; let $\theta$ be the trainable parameters of $v^{\mathrm{ft}}$.
3:for epoch $1,\dots$ do
4:  Sample $x_{0}\!\sim\!\mathcal{N}(0,I)$ and $\alpha_{0}\!\sim\!\mathcal{N}(0,I)$; $\bar{T}\leftarrow\textsc{get\_tilted\_time}(T)$.
5:
6:  for $i=0$ to L1L-1 do
7:   $\leftarrow\!t_{i}, t+t$$^{+}\!\leftarrow\!t_{i+1}$, $h\!\leftarrow\!t^{+}-t$.
8:   Base state velocity:
9:     $\mathbf{v}_{x,t}^{\mathrm{base}}=v_{x}^{\mathrm{base}}(x_{t},t)$.
10:   Surrogate terminal parameter:
11:     $\hat{x}_{1}=x_{t}^{\mathrm{ft}}+(1-t)\,\mathbf{v}_{x,t}^{\mathrm{base}}$
12:     $\hat{\alpha}_{1}^{\mathrm{ft}}=\varphi(\hat{x}_{1})$.
13:   Base parameter velocity (surrogate):
14:     $\mathbf{v}_{\alpha,t}^{\mathrm{base}}=(\hat{\alpha}_{1}^{\mathrm{ft}}-\alpha_{t})/(1-t)$.
15:   Fine-tuned velocities:
16:     $\mathbf{v}_{x,t}^{\mathrm{ft}}=v_{x}^{\mathrm{ft}}(x_{t},\alpha_{t},\mathbf{v}_{\alpha,t}^{\mathrm{base}},t)$
17:     $\mathbf{v}_{\alpha,t}^{\mathrm{ft}}=v_{\alpha}^{\mathrm{ft}}(x_{t},\alpha_{t},\mathbf{v}_{\alpha,t}^{\mathrm{base}},t)$.
18:   Forward steps (with noise $\sigma(t)$):
19:     $x_{t^{+}}^{\mathrm{base}}=\textsc{Step}(x_{t}^{\mathrm{base}},\mathbf{v}_{x,t}^{\mathrm{base}},\sigma(t),h)$
20:     $x_{t^{+}}^{\mathrm{ft}}=\textsc{Step}(x_{t}^{\mathrm{ft}},\mathbf{v}_{x,t}^{\mathrm{ft}},\sigma(t),h)$
21:     $\alpha_{t^{+}}^{\mathrm{ft}}=\textsc{Step}(\alpha_{t}^{\mathrm{ft}},\mathbf{v}_{\alpha,t}^{\mathrm{ft}},\sigma(t),h)$.
22:  end for
23:
24:  $(a_{x,t},a_{\alpha,t})_{t\in\bar{T}}\leftarrow\textsc{SolveLeanAdjoint}(\{x_{t}^{\mathrm{base}}\},\{x_{t}^{\mathrm{ft}}\},\{\alpha_{t}^{\mathrm{ft}}\})$.
25:
26:  Subsample time indices: form $\mathcal{I}\subset\{0,\dots,L\}$ according to $K$ and $L_{\text{last}}$
27:
28:  Per-step drifts: for each $t_{i}\in\mathcal{I}$ compute
29:     $\mathbf{b}_{x,t_{i}}^{\mathrm{ft}}=\mathbf{v}_{x,t_{i}}^{\mathrm{ft}}+\frac{\sigma(t_{i})^{2}}{2\,\eta(t_{i})}\bigl(\mathbf{v}_{x,t_{i}}^{\mathrm{ft}}-\frac{1}{t_{i}+\varepsilon}\,x_{t_{i}}^{\mathrm{ft}}\bigr)$;  $\mathbf{b}_{x,t_{i}}^{\mathrm{base}}=\mathbf{v}_{x,t_{i}}^{\mathrm{base}}+\frac{\sigma(t_{i})^{2}}{2\,\eta(t_{i})}\bigl(\mathbf{v}_{x,t_{i}}^{\mathrm{base}}-\frac{1}{t_{i}+\varepsilon}\,x_{t_{i}}^{\mathrm{base}}\bigr)$.
30:     $\mathbf{b}_{\alpha,t_{i}}^{\mathrm{ft}}=\mathbf{v}_{\alpha,t_{i}}^{\mathrm{ft}}+\frac{\sigma(t_{i})^{2}}{2\,\eta(t_{i})}\bigl(\mathbf{v}_{\alpha,t_{i}}^{\mathrm{ft}}-\frac{1}{t_{i}+\varepsilon}\,\alpha_{t_{i}}^{\mathrm{ft}}\bigr)$;  $\mathbf{b}_{\alpha,t_{i}}^{\mathrm{base}}=\mathbf{v}_{\alpha,t_{i}}^{\mathrm{base}}+\frac{\sigma(t_{i})^{2}}{2\,\eta(t_{i})}\bigl(\mathbf{v}_{\alpha,t_{i}}^{\mathrm{base}}-\frac{1}{t_{i}+\varepsilon}\,\alpha_{t_{i}}^{\mathrm{ft}}\bigr)$.
31:
32:  Per-step losses (with clipping): for each $t_{i}\in\mathcal{I}$ set
33:     $\ell_{x}(t_{i})=\min\!\bigl\{\,[\sigma(t_{i})^{-1}(\mathbf{b}_{x,t_{i}}^{\mathrm{ft}}-\mathbf{b}_{x,t_{i}}^{\mathrm{base}})+\sigma(t_{i})\,a_{x,t_{i}}]^{2},\;\mathrm{LCT}_{x}\,\bigr\}$,
34:     $\ell_{\alpha}(t_{i})=\min\!\bigl\{\,[\sigma(t_{i})^{-1}(\mathbf{b}_{\alpha,t_{i}}^{\mathrm{ft}}-\mathbf{b}_{\alpha,t_{i}}^{\mathrm{base}})+\sigma(t_{i})\,a_{\alpha,t_{i}}]^{2},\;\mathrm{LCT}_{\alpha}\,\bigr\}$,
35:     (squaring and clipping applied elementwise, then summed over spatial dimensions)
36:
37:  Adjoint–matching objective:
38:    $\mathcal{L}=\frac{1}{|\mathcal{I}|}\sum_{t_{i}\in\mathcal{I}}\bigl(\ell_{x}(t_{i})+\ell_{\alpha}(t_{i})\bigr)$.
39:    $\theta\leftarrow\textsc{gradient\_descent}(\theta,\mathcal{L})$
40:end for
41:return $v^{\mathrm{ft}},\;\varphi$
Appendix E Details: Experiments
Here we describe the fine-tuning configuration of the conducted experiments. Specifics about the base model FM training can be found in Appendix C. In all experiments, we use AdamW as the optimizer with a weight decay of $0.01$ and all fine-tuning is conducted on a single NVIDIA L40S GPU. As reported in the main paper, fine-tuning for Darcy takes less than one minute per epoch. For the other, computationally more expensive experiments, one epoch of fine-tuning takes around one minute (1-2 hours for 100 epochs).
E.1 Metrics
SSIM diversity.
In panel (a) of Figure 3, we report a SSIM-based (Wang et al., 2004) metric for diversity, which is implemented as follows:
Given a batch $\{\alpha_{i}\}_{i=1}^{B}$ of single-channel parameter maps scaled to $[0,1]$, we define diversity as the mean complement of the pairwise structural similarity index (SSIM) :
$\mathcal{D}_{\text{SSIM}}(\{\alpha_{i}\}_{i=1}^{B})\;=\;\frac{1}{\binom{B}{2}}\sum_{1\leq i<j\leq B}\Bigl(1-\operatorname{SSIM}(\alpha_{i},\alpha_{j})\Bigr),\qquad\alpha_{i}\in[0,1]^{H\times W}.$
With $\operatorname{SSIM}\in[0,1]$ (data range =1=1), $\mathcal{D}_{\text{SSIM}}\in[0,1]$ and larger values indicate greater sample diversity.
Maximum Mean Discrepancy (MMD).
To quantify distributional differences between generated samples and the reference dataset $\mathcal{D}_{\mathrm{ref}}$, we use the squared Maximum Mean Discrepancy (MMD2) with an RBF kernel (Gretton et al., 2012). Given two batches $\{x_{i}\}_{i=1}^{N_{1}}\subset\mathbb{R}^{D}$ and $\{y_{j}\}_{j=1}^{N_{2}}\subset\mathbb{R}^{D}$, let
$k_{\sigma}(u,v)=\exp\!\Bigl(-\tfrac{1}{2\sigma^{2}}\|u-v\|_{2}^{2}\Bigr),\qquad\gamma=\tfrac{1}{2\sigma^{2}}.$
The empirical MMD2 is
$\operatorname{MMD}^{2}(x,y)=\frac{1}{N_{1}^{2}}\sum_{i,i^{\prime}}k_{\sigma}(x_{i},x_{i^{\prime}})+\frac{1}{N_{2}^{2}}\sum_{j,j^{\prime}}k_{\sigma}(y_{j},y_{j^{\prime}})-\frac{2}{N_{1}N_{2}}\sum_{i,j}k_{\sigma}(x_{i},y_{j}).$
In all experiments we use the median heuristic (Gretton et al., 2012), which sets the kernel bandwidth $\sigma^{2}$ to the median of the pairwise squared distances across the pooled samples $z=\{x_{i}\}\cup\{y_{j}\}$. In practice we subsample up to $1000$ points to compute
$\sigma^{2}=\operatorname{median}\bigl\{\|z_{a}-z_{b}\|_{2}^{2}:a<b\bigr\}+\varepsilon,\qquad\varepsilon=10^{-6},$
which yields a stable and scale-adaptive kernel width (Garreau et al., 2017).
E.2 Ablations and Baselines
Single Adjoint Matching.
To isolate the effect of modeling a joint flow over $(x,\alpha)$, we include two ablations based on the vanilla Adjoint Matching (AM) objective of Domingo-Enrich et al. (2025). Both variants optimize the standard lean-adjoint control loss over the state trajectory $\{x_{t}\}$, using the same noise schedule, ODE discretization, and step subsampling as our full method.
Base AM freezes the inverse predictor $\varphi$. It is used only to compute PDE residuals but receives no gradient updates, and no joint flow over $\alpha$ is modeled.